heres the problem:

4(3x*4x)=18
my teacher gave the problem off the top of his head but i have done the check to it and it does'nt make sense...um can u help !

Answer by stanbon(48569)

Divide both sides by "4" to get 12x^2=9/2
Divide both sides by "12" to get x^2=3/8
x=sqrt3/2sqrt2
Multiply numerator anddenominator by sqrt2 to get:
x=[sqrt6]/4
Cheers,
Stan H.